OpenStudy (anonymous):

You are to manufacture a rectangular box with 3 dimensions x, y and z, and volume v=729. Find the dimensions

OpenStudy (anonymous):

Any other restrictions? Could just have x=y=z=cube root of 729..

OpenStudy (anonymous):

NNO THAT ALL INF THAT I HAVE

OpenStudy (anonymous):

It's pretty open-ended then. There are an infinite number of solutions. x=y=z=9 is the simplest, but you can pick any values you want for 2 of the variables then solve for the third and it would be valid.

OpenStudy (anonymous):

i dont know coz that the way the qustion was written and it asking to solve x ? y ? z? and i dont know how to do it

OpenStudy (anonymous):

Start with the rectangular prism volume formula, V=x*y*z V = 729, which is a perfect cube (729=9^3), so let all three dimensions =9 and that will satisfy the conditions. If you want something other than a cube, then pick any numbers you want for, say, x&y: e.g. x=4, y=10, then 729 = 4*10*z, so z=729/40 = 18.225. If you're not restricted to whole numbers or even rational numbers, then you can use whatever you want.